# Complexity of Linear Subsequences of $k$-Automatic Sequences

### v2

@@ Abstract
We construct automata with input(s) in base $k$ recognizing some basic relations and study their number of states. We also consider some basic operations on $k$-automatic sequences and discuss their state complexity. We find a relationship between subword complexity of the interior sequence $(h^{\prime}(i))_{i\geq 0}$ and state complexity of the linear subsequence $(h(ni+c))_{i\geq 0}$. We resolve a recent question of Zantema and Bosma about linear subsequences of $k$-automatic sequences with input in most-significant-digit-first format. We also discuss the state complexity and runtime complexity of using a reasonable interpretation of Büchi arithmetic to actually construct some of the studied automata recognizing relations or carrying out operations on automatic sequences.
@@ 1 Introduction
In this paper, we are concerned with arithmetic in base $k$ and with $k$-automatic sequences for integers $k\geq 2$. This interesting class of sequences was introduced by Cobham [10, 11] and has been studied for more than 50 years; see [2] for a monograph on automatic sequences. A sequence $(h(i))_{i\geq 0}$ is said to be $k$-automatic if there is a deterministic finite automaton with output (DFAO) computing the $i$-th term, given the base-$k$ representation of $i$ as input. A DFAO is a $6$-tuple $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ where
@@
We say a DFAO $M$ generates the sequence ${\bf h}=(h(i))_{i\geq 0}$ if $\tau(\delta(q_{0},x))=h(i)$ for all $i\geq 0$ and all words $x\in\Sigma_{k}^{*}$ such that $x]_{k}=i$. In particular, $M$ is expected to give the correct result no matter how many leading zeros are appended to the input. The state complexity of an automatic sequence is defined to be the number of states in a minimal DFAO generating it. A DFAO with base-$k$ input is called a $k$-DFAO.
For a given automatic sequence ${\bf h}=(h(i))_{i\geq 0}$ there is a unique associated interior sequence ${\bf h}^{\prime}=(h^{\prime}(i))_{i\geq 0}$ with output values from $Q$, defined by taking the minimal DFAO $(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ and replacing $\tau$ by the identity map $i$ on $Q$. Here “unique” means up to renaming of the letters. It follows that $(h(i))_{i\geq 0}$ is the image of $(h^{\prime}(i))_{i\geq 0}$ under the coding $q\rightarrow\tau(q)$.
For a DFAO $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$, let the DFAO $\hat{M}=(Q,\Sigma_{k},\delta,q_{0},Q,i)$ where $i:Q\rightarrow Q$ is the identity map be the intrinsic DFAO of $M$ and let $\hat{\bf M}$ be the sequence generated by the intrinsic DFAO $\hat{M}$. The interior sequence $(h^{\prime}(i))_{i\geq 0}$ of an automatic sequence $(h(i))_{i\geq 0}$ is the sequence generated by the intrinsic DFAO of the minimal DFAO for $(h(i))_{i\geq 0}$.
A classical result of Cobham [11, p. 174] states that if $(h(i))_{i\geq 0}$ is $k$-automatic, then so is the linear subsequence $(h(ni+c))_{i\geq 0}$ for integers $n\geq 1$ and $c\geq 0$. In this paper, we are interested in the state complexity of taking linear subsequences of automatic sequences and related topics. Our results on the state complexity of linear subsequences expand on the paper by Zantema and Bosma [23] and answer one of their open questions. More details can be found in Section 4.
@@
Boudet and Comon [6] and Wolper and Boigelot [22] also discussed the size of the automaton for an arbitrary formula in Presburger arithmetic. Another motivation for studying the topics in this paper is their appearance in expressions in Büchi arithmetic, which has been implemented (for example) in the Walnut software system [18] for solving problems in combinatorics on words. Users of this system want to know how long some of the basic expressions can take to evaluate, and how big the resulting automata are. In this paper, we consider the automata for specific relations and automatic sequences. Here we have to examine different considerations than simply state complexity (which measures the size of the smallest possible finite automaton for a language or sequence), because we need to know how large the intermediate automata can be in the constructions implied by Büchi arithmetic. In particular, we need to consider (a) how an expression in Büchi arithmetic is translated to an automaton, (b) what are the sizes of the intermediate automata created in this translation, and (c) what the total running time of the procedure is.
In Section 2, we introduce the required background. The rest of the paper is dedicated to discussing our results and their relevance to previously known results. In Section 3, we discuss results regarding the state complexity of recognizing additive and multiplicative relations with base-$k$ input. In Section 4, we study the state complexity of linear subsequences of a $k$-automatic sequence, and find a relationship between this state complexity and subword complexity. In Section 5, we analyze the runtime complexity of constructing some of the automata studied in previous sections using an interpretation of Büchi arithmetic. Finally, in Section 6, we state some open problems.
@@ 2 Background
The canonical base-$k$ representation of an integer $i$ is the word $x$, without leading zeros, such that $x]_{k}=i$. Furthermore, we use the notation $(i)_{k}$ for the canonical representation of $i$ in base $k$. The canonical representation of 0 is $\lambda$, the empty word. Sometimes we use $[x]$ instead of $x]_{k}$ and $(i)$ instead of $(i)_{k}$. We say a word $z$ is a valid right extension of a word $x$ if $x$ is a prefix of $z$.
@@
Finally, we also need the notion of nondeterministic finite automaton (NFA), which is like that of a DFA, except that the transition function $\delta:2^{Q}\times\Sigma\rightarrow 2^{Q}$ is now set-valued. Here an input is accepted if $\delta(q_{0},x)$ contains some element in $A$.
@@
Let $\bf x$} be an automatic sequence generated by an $m$-state DFAO $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ with msd-first input and let $\hat{\bf M}$ be the sequence generated by the intrinsic DFAO $\hat{M}=(Q,\Sigma,\delta,q_{0},Q,i)$, where $i$ is the identity map. Then $\rho_{\bf x}(n)\leq kn\rho_{\hat{\bf M}}(2)\leq knm^{2}$ for all $n\geq 1$.
@@
See [2, Theorem 10.3.1]. We point out that the statement of the theorem in [2] only gives $\rho_{\bf x}(n)\leq knm^{2}$, but the proof actually presented there provides the other inequality. ∎
@@ 3 Recognizing Relations
We say a DFA recognizes a relation (such as the addition relation $i+j=l$) if it takes the values of the variables in parallel, represented in some base $k$, and accepts if and only if the relation holds. In this section we review some of the basic arithmetic operations and what is known about their state complexity.
@@ 3.1 Addition
Now we consider what happens when we specify one summand.
@@
On input $x\times y$, we need to keep track of the difference $D_{k}(x,y)=[y]_{k}-[x]_{k}$ and accept the input if and only if the difference is $c$. Each state of the automaton is a possible difference $d$, but we need to find a range $I$ for $d$ so that the automaton is actually finite. We have
@@
Therefore, if $D(x,y)$ is outside the range $[0,c]$ it will not return to it for any right extension of $x\times y$, and we can take $I$ to be the interval $[0,c]$. The transition function is based on Eq. (1). After reading each input letter pair, the difference either stays the same or increases (unless it is no longer in the range $I$). Furthermore, if we consider a state $d\neq 0$, there are only two other states that have a direct transition to $d$, namely $\lfloor\frac{d}{k}\rfloor$ and $\lceil\frac{d}{k}\rceil$.
So we construct the automaton $M_{c}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ where $\delta(d,[a,b])=kd+b-a$, $q_{0}=0$, and $A=\{c\}$. To create the set $Q$, we start with $\{c\}$. Then, we add the states $c^{\prime}=\lfloor\frac{c}{k}\rfloor$ and $c^{\prime\prime}=\lceil\frac{c}{k}\rceil$ to $Q$. We know $|c^{\prime\prime}-c^{\prime}|\leq 1$; therefore, the set $\{\lfloor\frac{c^{\prime}}{k}\rfloor,\lfloor\frac{c^{\prime\prime}}{k}\rfloor,\lceil\frac{c^{\prime}}{k}\rceil,\lceil\frac{c^{\prime\prime}}{k}\rceil\}$ has size at most $2$. We repeat the same step (adding the floor and ceiling of division by $k$ to $Q$) until reaching the state 0. This step is repeated $O(\log_{k}c)$ times and there are always at most two new states created at each step. Therefore, the number of states in $M_{c}$ is $O(\log_{k}c)$.
@@
$\delta([l,r],[a,b])$ $=\begin{cases}[l,(r+a-b)/k],&\text{if $l=\lfloor\log_{k}c+1\rfloor$};\\ [l+1,(r+c_{l}+a-b)/k],&\text{otherwise,}\end{cases}$
@@
Consider the construction of automaton $M_{c}$ for msd-first input described in the proof of Theorem 4. Let us group the new states created at each step of the construction into a level including an initial level for the state $c$. We call a level with $1$ state a single-state level and a level with $2$ states a double-state level.
@@
However, for some values of $k$ and $c$, some double-state levels in the simple form are changed to single-state levels causing the number of states to be lower than the simple form. We study these cases, how they are reflected in the theorem’s formula, and why they are the only possible such cases.
@@
If $k$ divides some integer $c^{\prime}$, then the floor and ceiling of $\frac{c^{\prime}}{k}$ is the same number. As a result, in this case, instead of the last level being single-state, the last $1+\nu_{k}(c)$ levels are single-state. So Case 1 is covered by the $\nu_{k}(c)$ in the theorem’s formula. The automaton $M_{6}$ in base $3$ is an example of this case.
@@
Consider a double-state level consisting of states $d$ and $d+1$ following a single-state level $d^{\prime}$. In this case, $d$ transits to $d+1$ (proof by contradiction). We have d+1=kd+bad+1=kd+b-a for some $a,b\in[0,k-1]$, $d\neq 0$, and $k\geq 2$. The only possible solutions to this equation are the following:
@@
If some positive integer $c^{\prime}$ is not divisible by $k$, then the floor and ceiling of $\frac{c^{\prime}}{k}$ are different numbers, except for the special case already discussed in Case 2. So Case 1 covers all the other single-state levels not counted in Case 2.
@@ 3.2 Multiplication
In addition to providing an msd-first automaton construction for recognizing the relation $x]_{k}+[y]_{k}=[z]_{k}$, Waxweiler [21] provided one for recognizing $n[x]_{k}=[y]_{k}$; and later, Charlier et al. [9] provided one for recognizing $n[x]_{k}+c=[y]_{k}$. Furthermore, the automaton for $n[x]_{k}+c=[y]_{k}$ with msd-first base $2$ and lsd-first base $2$ input can be created by the methods explained respectively by Wolper and Boigelot [22] and Boudet and Comon [6]. The following theorem considers the automaton accepting $n[x]_{k}+c=[y]_{k}$ in the case of $c<n$ with both lsd- and msd-first base $k$ input.
@@
However, it is easy to prove that
@@
4 Linear Subsequences
Zantema and Bosma [23] have shown that for lsd-first input, if $(h(i))_{i\geq 0}$ is a $k$-automatic sequence generated by a DFAO of $m$ states, then $(h(i+1))_{i\geq 0}$ is generated by a DFAO of at most $2m$ states. Furthermore, they have shown that for msd-first input, $(h(i+1))_{i\geq 0}$ is generated by a DFAO of at most $m^{2}$ states and there is an example where this bound is tight. Additionally, they have shown that the same bounds hold for $(h(i-1))_{i\geq 0}$ where $h(0)$ is any chosen letter from sequence alphabet.
However, they did not show that their bound of $2m$ states for $(h(i+1))_{i\geq 0}$ in the lsd-first case is tight. We suspect it is not. In the following result, we show there are examples attaining the very slightly weaker bound of 2m12m-1 states.
@@ 4 Linear Subsequences
For each $m\geq 2$, there exists a $m$-state DFAO $M$ in lsd-first format generating the sequence $(h(i))_{i\geq 0}$ such that at least 2m12m-1 states are required in the lsd-first DFAO $M^{\prime}$ generating $(h(i+1))_{i\geq 0}$.
@@
We use an adaptation of the Myhill-Nerode theorem [14, §3.4] for DFAOs. The adaptation for DFAOs is also used by Zantema and Bosma [23].
@@
$\delta(q,a)$ $=\begin{cases}(q+1)\bmod m,&\text{if $a=1$},\\ q+1,&\text{if $a=0$ and $q\in\{0,\ldots,m-3\}$},\\ q,&\text{if $a=0$ and $q\in\{m-2,m-1\}$;}\end{cases}$
@@
$\tau(q)$ $=\begin{cases}1,&\text{if $q=m-1$},\\ 0,&\text{otherwise}.\end{cases}$
@@
Case 1: $0^{i}$ and $0^{j}$ where $1\leq i<j\leq m-2$. Concatenate $0^{m-j-2}1$. The new words are $0^{i+m-j-2}1$ and $0^{m-2}1$. The output of automaton $M^{\prime}$ on these words is then the same as the output of automaton $M$ on inputs $10^{i-j+m-3}1$ and $10^{m-3}1$, which are 0 and $1$. Therefore, the two words are distinguishable.
@@
Let the DFAO generating the sequence $(h(i))_{i\geq 0}$ be $M=(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$. We want to design a DFAO $M^{\prime}$ for $(h(i+j))_{i,j\geq 0}$. The basic idea is to perform $i+j$ while reading the representation of $i$ and $j$ in parallel; however, upon reading a letter pair $[a,b]$ and before reading the next letter pair, we do not know what the letter corresponding to $[a,b]$ in $i+j$ is. For example, $00]_{3}+[01]_{3}=[01]_{3}$ but $01]_{3}+[02]_{3}=[10]_{3}$; the first letter pairs read are the same but the first letters in the addition results are not. So each state in automaton $M^{\prime}$ includes two states from $M$: one of them corresponds to the case where a carry is not added to $a+b$, and one to the case where a carry is added.
More formally, we construct automaton $M^{\prime}=(Q^{\prime},\Sigma_{k}^{2},\delta^{\prime},q^{\prime}_{0},\Delta,\tau^{\prime})$ as follows.
@@
$\delta^{\prime}([q,p],[a,b])$ $=\begin{cases}[\delta(q,a+b),\delta(q,a+b+1)],&\text{if $a+b\leq k-2$},\\ [\delta(q,a+b),\delta(p,a+b+1-k)],&\text{if $a+b=k-1$},\\ [\delta(p,a+b-k),\delta(p,a+b+1-k)],&\text{if $a+b\geq k$;}\end{cases}$
@@
The automaton is constructed so that if $[q,p]$ is a state in automaton reachable by $x\times y$, then $q,p]=[h^{\prime}([x]_{k}+[y]_{k}),h^{\prime}([x]_{k}+[y]_{k}+1)]$. So the reachable states are length-$2$ subsequences of the interior sequence $(h^{\prime}(i))_{i\geq 0}$. Therefore, based on Theorem 2, there are $O(m^{2})$ reachable states. ∎
@@
In the special case that $0\leq c<n$, this lsd-first bound was improved slightly to at most $mn$ states by Zantema and Bosma [23]. Furthermore, they provided examples where the same bound does not hold for msd-first input. They stated that obtaining a tight bound in the msd-first case as an open question. We address this in the following theorem, which provides a close connection between state complexity and subword complexity.
@@
    If $c<n$, then there is a DFAO of at most $\rho_{{\bf h}^{\prime}}(n)$ states generating $(h(ni+c))_{i\geq 0}$.
@@
    If $c\geq n$, then there is a DFAO of at most $\rho_{{\bf h}^{\prime}}(c+1)$ states generating $(h(ni+c))_{i\geq 0}$.
@@
    Suppose $c<n$. We want to create a DFAO generating $(h(ni+c))_{i\geq 0}$.
    The idea is that we construct $M^{\prime}$ such that on input $x$ it reaches a state given by the following subword of the interior sequence $(h^{\prime}(i))_{i\geq 0}$:
@@
    If $c<n$, then there is a DFAO of at most $knm^{2}$ states with msd-first input generating the automatic sequence $(h(ni+c))_{i\geq 0}$.
@@
    If $c\geq n$, then there is a DFAO of at most $k(c+1)m^{2}$ states with msd-first input generating the automatic sequence $(h(ni+c))_{i\geq 0}$.

### v3

@@ Abstract
We construct automata with input(s) in base $k$ recognizing some basic relations and study their number of states. We also consider some basic operations on $k$-automatic sequences $(h(i))_{i\geq 0}$ and discuss their state complexity. We find a relationship between subword complexity of the interior sequence $(h^{\prime}(i))_{i\geq 0}$ and state complexity of the linear subsequence $(h(ni+c))_{i\geq 0}$. We resolve a recent question of Zantema and Bosma about linear subsequences of $k$-automatic sequences with input in most-significant-digit-first format. We also discuss the state complexity and runtime complexity of using a reasonable interpretation of Büchi arithmetic to actually construct some of the studied automata recognizing relations or carrying out operations on automatic sequences.
@@ 1 Introduction
Determining the number of states in a minimal automaton for a formal language or automatic sequence is an extensively studied area of formal languages and automata theory called state complexity. The reader can refer to works by Yu et al. [29], Gao et al. [14], and Zantema and Bosma [30] for some examples. This paper includes discussions and results on the state complexity of specific languages and automatic sequences with input in base $k$.
The $k$-automatic sequences for integers $k\geq 2$ is an interesting class of sequences introduced by Cobham [12, 13] and was studied for more than 50 years; see [2] for a monograph on automatic sequences. A sequence $(h(i))_{i\geq 0}$ is said to be $k$-automatic if there is a deterministic finite automaton with output (DFAO) computing the $i$-th term, given the base-$k$ representation of $i$ as input. A DFAO is a $6$-tuple $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ where
@@
We say a DFAO $M$ generates the sequence ${\bf h}=(h(i))_{i\geq 0}$ if $\tau(\delta(q_{0},x))=h(i)$ for all $i\geq 0$ and all words $x\in\Sigma_{k}^{*}$ such that $x]_{k}=i$. In particular, $M$ is expected to give the correct result no matter how many leading zeros are appended to the input. A DFAO with base-$k$ input is called a $k$-DFAO.
For a given automatic sequence ${\bf h}=(h(i))_{i\geq 0}$ there is a unique associated interior sequence ${\bf h}^{\prime}=(h^{\prime}(i))_{i\geq 0}$ with output values from $Q$, defined by taking the minimal DFAO $(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ and replacing $\tau$ by the identity map $\iota$ on $Q$. Here “unique” means up to renaming of the letters. It follows that $(h(i))_{i\geq 0}$ is the image of $(h^{\prime}(i))_{i\geq 0}$ under the coding $q\rightarrow\tau(q)$.
For a DFAO $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$, let the DFAO $\hat{M}=(Q,\Sigma_{k},\delta,q_{0},Q,\iota)$, where $\iota:Q\rightarrow Q$ is the identity map, be the intrinsic DFAO of $M$ and let $\hat{\bf M}$ be the sequence generated by the intrinsic DFAO $\hat{M}$. The interior sequence $(h^{\prime}(i))_{i\geq 0}$ of an automatic sequence $(h(i))_{i\geq 0}$ is the sequence generated by the intrinsic DFAO of the minimal DFAO for $(h(i))_{i\geq 0}$.
A classical result of Cobham [13, p. 174] states that if $(h(i))_{i\geq 0}$ is $k$-automatic, then so is the linear subsequence $(h(ni+c))_{i\geq 0}$ for integers $n\geq 1$ and $c\geq 0$. In this paper, we are interested in the state complexity of linear subsequences of automatic sequences and related topics. Our results on the state complexity of linear subsequences expand on the paper by Zantema and Bosma [30] and answer one of their open questions. More details can be found in Section 4.
@@
At first glance, it might seem that results for the state complexity of problems involving msd-first and lsd-first representations would be closely related. While this is true for arithmetic, it is not true for subsequences of automatic sequences. Roughly speaking, this is because a finite-state transducer can carry out addition and multiplication by constants on lsd-first representations, but not msd-first representations.
Boudet and Comon [7] and Wolper and Boigelot [28] also discussed the size of the automaton for an arbitrary formula in Presburger arithmetic. Furthermore, Amat et al. [3] have developed a tool named Patapsco that computes a lower bound for the number of states in the automaton for a given Presburger arithmetic formula. This tool computes a lower bound without actually constructing the automaton.
Another motivation for studying the topics in this paper is their appearance in expressions in Büchi arithmetic, which was implemented (for example) in the Walnut software system [24] for solving problems in combinatorics on words. Users of this system want to know how long some of the basic expressions can take to evaluate, and how big the resulting automata are. In this paper, we consider the automata for specific relations and automatic sequences. Here we have to examine different considerations than simply state complexity (which measures the size of the smallest possible finite automaton for a language or sequence), because we need to know how large the intermediate automata can be in the constructions implied by Büchi arithmetic. In particular, we need to consider (a) how an expression in Büchi arithmetic is translated to an automaton, (b) what are the sizes of the intermediate automata created in this translation, and (c) what the total running time of the procedure is.
In Section 2, we introduce the required background. The rest of the paper is dedicated to discussing our results and their relevance to previously known results. In Section 3, we discuss results regarding the state complexity of recognizing additive and multiplicative relations with base-$k$ input. In Section 4, we study the state complexity of linear subsequences of a $k$-automatic sequence, and find a novel relationship between this state complexity and subword complexity. We also solve an open problem of Zantema and Bosma. This section contains our main novel results. In Section 5, we analyze the runtime complexity of constructing some of the automata studied in previous sections using an interpretation of Büchi arithmetic. Finally, in Section 6, we state some open problems.
@@ 2 Background
The canonical base-$k$ representation of an integer $i$ is the word $x$, without leading zeros, such that $x]_{k}=i$. Furthermore, we use the notation $(i)_{k}$ for the canonical representation of $i$ in base $k$. The canonical representation of 0 is $\lambda$, the empty word. Sometimes, if the context is clear, we just use $[x]$ instead of $x]_{k}$ and $(i)$ instead of $(i)_{k}$. We say a word $z$ is a valid right extension of a word $x$ if $x$ is a prefix of $z$.
@@
Finally, we also need the notion of nondeterministic finite automaton (NFA), which is like that of a DFA, except that the transition function $\delta:Q\times\Sigma\rightarrow 2^{Q}$ has set-valued output. Furthermore, the transition function can be extended to domain $2^{Q}\times\Sigma$. Here an input is accepted if $\delta(q_{0},x)$ contains some element in $A$.
The state complexity of a formal language is the number of states in the minimal DFA recognizing it and we say the state complexity of an automatic sequence is the number of states in the minimal DFAO generating it.
@@
Let $\bf x$} be an automatic sequence generated by an $m$-state DFAO $M=(Q,\Sigma,\delta,q_{0},\Delta,\tau)$ with msd-first input and let $\hat{\bf M}$ be the sequence generated by the intrinsic DFAO $\hat{M}=(Q,\Sigma,\delta,q_{0},Q,\iota)$, where $\iota$ is the identity map. Then $\rho_{\bf x}(n)\leq kn\rho_{\hat{\bf M}}(2)\leq knm^{2}$ for all $n\geq 1$.
@@
See [2, Theorem 10.3.1]. We point out that the statement of the theorem in [2] only gives the inequality $\rho_{\bf x}(n)\leq knm^{2}$, but the proof actually presented there provides the other inequality. ∎
@@ 3 Recognizing Relations
We say a DFA recognizes a relation (such as the addition relation $i+j=l$) if it takes the values of the variables in parallel, represented in some base $k$, and accepts if and only if the relation holds. In this section we review some of the basic arithmetic operations, what is known about their state complexity, and some new results.
@@ 3.1 Addition
Now we consider what happens when we specify one summand to be a fixed constant.
@@
On input $x\times y$, we need to keep track of the difference $D_{k}(x,y)=[y]_{k}-[x]_{k}$ and accept the input if and only if the difference is $c$. Each state of the automaton is a possible difference $d$, but we need to find a suitable range $I$ for $d$ so that the automaton is actually finite. We have
@@
Therefore, if $D(x,y)$ is outside the range $[0,c]$ it will not return to it for any right extension of $x\times y$, and we can take $I$ to be the interval $[0,c]$. The transition function is based on Eq. (1). After reading each input letter pair, the difference either stays the same or increases (unless it is no longer in the range $I$). Furthermore, if we consider a state $d\neq 0$, there are only two other states that have a direct transition to $d$, namely $\lfloor d/k\rfloor$ and $\lceil d/k\rceil$.
So we construct the automaton $M_{c}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ where $\delta(d,[a,b])=kd+b-a$, $q_{0}=0$, and $A=\{c\}$. To create the set $Q$, we start with $\{c\}$. Then we add the states $c^{\prime}=\lfloor c/k\rfloor$ and $c^{\prime\prime}=\lceil c/k\rceil$ to $Q$. We know $|c^{\prime\prime}-c^{\prime}|\leq 1$; therefore, the set $\{\lfloor c^{\prime}/k\rfloor,\lfloor c^{\prime\prime}/k\rfloor,\lceil c^{\prime}/k\rceil,\lceil c^{\prime\prime}/k\rceil\}$ has size at most $2$. We repeat the same step (adding the floor and ceiling of division by $k$ to $Q$) until reaching the state 0. This step is repeated $O(\log_{k}c)$ times and there are always at most two new states created at each step. Therefore, the number of states in $M_{c}$ is $O(\log_{k}c)$.
@@
$\delta([l,r],[a,b])$ $=\begin{cases}[l,(r+a-b)/k],&\text{if $l=\lfloor\log_{k}c+1\rfloor$};\\ [l+1,(r+c_{l}+a-b)/k],&\text{otherwise;}\end{cases}$
@@
Consider the construction of the automaton $M_{c}$ for msd-first input described in the proof of Theorem 4. Let us group the new states created at each step of the construction into a level including an initial level for the state $c$. We call a level with $1$ state a single-state level and a level with $2$ states a double-state level.
@@
However, for some values of $k$ and $c$, some double-state levels in the simple form are changed to single-state levels, causing the number of states to be lower than the simple form. We study these cases, how they are reflected in the theorem’s formula, and why they are the only possible such cases.
@@
If $k$ divides some integer $c^{\prime}$, then $\lfloor c^{\prime}/k\rfloor=\lceil c^{\prime}/k\rceil$. As a result, in this case, instead of the last level being single-state, the last $\nu_{k}(c)+1$ levels are single-state. So Case 1 is covered by the $\nu_{k}(c)$ in the theorem’s formula. The automaton $M_{6}$ in base $3$ is an example of this case.
@@
Consider a double-state level consisting of states $d$ and $d+1$ following a single-state level $d^{\prime}$. In this case, $d$ transitions to $d+1$ (proof by contradiction). We have d+1=kd+bad+1=kd+b-a for some $a,b\in[0,k-1]$, $d\neq 0$, and $k\geq 2$. The only possible solutions to this equation are the following:
@@
If some positive integer $c^{\prime}$ is not divisible by $k$, then the floor and ceiling of $c^{\prime}/k$ are different numbers, except for the special case already discussed in Case 2. So Case 1 covers all the other single-state levels not counted in Case 2.
@@ 3.2 Multiplication
In addition to providing an msd-first automaton construction for recognizing the relation $x]_{k}+[y]_{k}=[z]_{k}$, Waxweiler [27] provided one for recognizing $n[x]_{k}=[y]_{k}$. Later, Charlier et al. [11] provided a construction for recognizing $n[x]_{k}+c=[y]_{k}$. Furthermore, the automaton for $n[x]_{k}+c=[y]_{k}$ with msd-first base-$2$ input (resp., lsd-first base-$2$ input) can be created by the methods explained by Wolper and Boigelot [28] (resp., Boudet and Comon [7]). The following theorem considers the automaton accepting $n[x]_{k}+c=[y]_{k}$ in the case $c<n$ with both lsd- and msd-first base $k$ input.
@@
However, it is easy to see that
@@
4 Linear Subsequences of Automatic Sequences
We now turn to examining the state complexity of transformations on automatic sequences.
Zantema and Bosma [30, Theorem 16] showed that for lsd-first input, if $(h(i))_{i\geq 0}$ is a $k$-automatic sequence generated by a DFAO of $m$ states, then $(h(i+1))_{i\geq 0}$ is generated by a DFAO of at most $2m$ states. Furthermore, in the same theorem they showed that for msd-first input, the sequence $(h(i+1))_{i\geq 0}$ is generated by a DFAO of at most $m^{2}$ states and there is an example where this bound is tight. Additionally, they showed [30, Theorem 17] that the same bounds hold for $(h(i-1))_{i\geq 0}$ where $h(0)$ is any chosen letter from the sequence alphabet.
However, they did not show that their upper bound of $2m$ states for $(h(i+1))_{i\geq 0}$ in the lsd-first case is tight. We suspect it is not. In the following result, we show there are examples attaining the very slightly weaker bound of 2m12m-1 states.
@@ 4 Linear Subsequences of Automatic Sequences
For lsd-first input, for each $m\geq 2$, there exists a $m$-state DFAO $M$ generating the sequence $(h(i))_{i\geq 0}$ such that at least 2m12m-1 states are required in the DFAO $M^{\prime}$ generating $(h(i+1))_{i\geq 0}$.
@@
We use an adaptation of the Myhill-Nerode theorem [17, §3.4] for DFAOs. The adaptation for DFAOs is also used by Zantema and Bosma [30] and the state equivalence for DFAOs is defined by van Spaendonck [25].
@@
$\delta(q,a)$ $=\begin{cases}(q+1)\bmod m,&\text{if $a=1$};\\ q+1,&\text{if $a=0$ and $q\in\{0,\ldots,m-3\}$};\\ q,&\text{if $a=0$ and $q\in\{m-2,m-1\}$;}\end{cases}$
@@
$\tau(q)$ $=\begin{cases}1,&\text{if $q=m-1$};\\ 0,&\text{otherwise}.\end{cases}$
@@
Case 1: $0^{i}$ and $0^{j}$ where $1\leq i<j\leq m-2$. Concatenate $0^{m-j-2}1$. The new words are $0^{i+m-j-2}1$ and $0^{m-2}1$. The output of the automaton $M^{\prime}$ on these words is then the same as the output of the automaton $M$ on inputs $10^{i-j+m-3}1$ and $10^{m-3}1$, which are 0 and $1$. Therefore, the two words are distinguishable.
@@
Let the DFAO generating the sequence $(h(i))_{i\geq 0}$ be $M=(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$. We want to design a DFAO $M^{\prime}$ for $(h(i+j))_{i,j\geq 0}$. The basic idea is to perform $i+j$ while reading the representation of $i$ and $j$ in parallel; however, upon reading a letter pair $[a,b]$ and before reading the next letter pair, we do not know what the letter corresponding to $[a,b]$ in $i+j$ is. For example, $00]_{3}+[01]_{3}=[01]_{3}$ but $01]_{3}+[02]_{3}=[10]_{3}$; the first letter pairs read are the same but the first letters in the addition results are not. So each state in the automaton $M^{\prime}$ includes two states from $M$: one of them corresponds to the case where a carry is not added to $a+b$, and one to the case where a carry is added.
More formally, we construct the automaton $M^{\prime}=(Q^{\prime},\Sigma_{k}^{2},\delta^{\prime},q^{\prime}_{0},\Delta,\tau^{\prime})$ as follows.
@@
$\delta^{\prime}([q,p],[a,b])$ $=\begin{cases}[\delta(q,a+b),\delta(q,a+b+1)],&\text{if $a+b\leq k-2$};\\ [\delta(q,a+b),\delta(p,a+b+1-k)],&\text{if $a+b=k-1$};\\ [\delta(p,a+b-k),\delta(p,a+b+1-k)],&\text{if $a+b\geq k$;}\end{cases}$
@@
The automaton is constructed so that if $[q,p]$ is a state in the automaton reachable by $x\times y$, then $q,p]=[h^{\prime}([x]_{k}+[y]_{k}),h^{\prime}([x]_{k}+[y]_{k}+1)]$. So the reachable states are length-$2$ subsequences of the interior sequence $(h^{\prime}(i))_{i\geq 0}$. Therefore, based on Theorem 2, there are $O(m^{2})$ reachable states. ∎
@@
In the special case that $0\leq c<n$, this lsd-first bound was improved slightly to at most $mn$ states by Zantema and Bosma [30]. Furthermore, they provided examples where the same bound does not hold for msd-first input. They stated obtaining a bound in the msd-first case as an open question. We address this in the following theorem, which provides a close and novel connection between state complexity and subword complexity.
@@
    If $c<n$, then there is a DFAO of at most $\rho_{{\bf h}^{\prime}}(n)$ states generating $(h(ni+c))_{i\geq 0}$ with msd-first input.
@@
    If $c\geq n$, then there is a DFAO of at most $\rho_{{\bf h}^{\prime}}(c+1)$ states generating $(h(ni+c))_{i\geq 0}$ with msd-first input.
@@
    Suppose $c<n$. We want to create a DFAO generating $(h(ni+c))_{i\geq 0}$. The idea is that we construct $M^{\prime}$ such that on input $x$ it reaches a state given by the following subword of the interior sequence $(h^{\prime}(i))_{i\geq 0}$:
@@
    If $c<n$, then there is a DFAO of at most $knm^{2}$ states generating the automatic sequence $(h(ni+c))_{i\geq 0}$ with msd-first input.
@@
    If $c\geq n$, then there is a DFAO of at most $k(c+1)m^{2}$ states generating the automatic sequence $(h(ni+c))_{i\geq 0}$ with msd-first input.